Marie has $10,000 to invest. She decided to borrow some funds at the risk-free rate of 6 percent to increase her investment in a portfolio with an expected return of 25 percent and a standard deviation of 30 percent. What is the expected return and standard deviation for her portfolio if she borrowed 50 percent of the portfolio value?

Expected Return?

Standard Deviation?
Answer #1

Amount borrowed =50%*10000 =5000
Total Investment =15000
Weight of Portfolio =15000/10000 =1.5
Weight of Risk free rate =-0.5
Expected Return =1.5*25%-0.5*6% =34.50%

Since Standard Deviation of Risk free asset =0
Hence the formula
Standard Deviation of Portfolio=Weight of Portfolio *Standard Deviation of Portfolio =1.5*30% =45%